OpenStudy (anonymous):

Simplify. 3 square root of 5 end root minus 2 square root of 2 end root plus square root of 45 end root minus square root of 28 2 square root of 12 2 square root of 2 6 square root of 5 end root minus 4 square root of 7 6 square root of 10 end root minus 4 square root of 14

OpenStudy (anonymous):

please someone? i can eliminate D

OpenStudy (mayankdevnani):

\[\large \bf 3 \sqrt{5}-2\sqrt{2}+\sqrt{45}-\sqrt{28}\] is this your question? @A.K.C.15

OpenStudy (anonymous):

the 2nd 2 is a 7

OpenStudy (mayankdevnani):

i can't understand your question

OpenStudy (mayankdevnani):

can you please write your question using LATEX.

OpenStudy (anonymous):

can i screen shot @mayankdevnani

OpenStudy (mayankdevnani):

yes,you can

OpenStudy (mayankdevnani):

first of all,you write wrong question,see there is no digit like 7 in your question

OpenStudy (anonymous):

oh oppss ugh @mayankdevnani

OpenStudy (anonymous):

thanks for making me notice

OpenStudy (mayankdevnani):

  so we get, \[\large \bf 3\sqrt{5}-2\sqrt{7}+3\sqrt{5}-2\sqrt{7}\] Take \(\large \bf \sqrt{5}~and~\sqrt{7}\) common,and then simplify,we get \[\large \bf \sqrt{5}(3+3)-\sqrt{7}(2+2)=6\sqrt{5}-4\sqrt{7}\]